Consider the reaction: 2H2 + O2 --> 2H2O. If a mixture of 1.0 mole of H2 and 1.0 mole O2 is taken, how many grams of H2O can form?

Consider the reaction: 3H2 + N2 --> 2NH3. If a mixture of 3.0 mole of H2 and 3.0 mole N2 is taken, how many moles of excess reactant remain unused?

The fermentation of glucose C6H12O6 produces ethanol, C2H5OH and CO2.The equation for the reaction is as follows: C6H12O6 --> 2C2H5OH + 2CO2.

How many moles of CO2 are produced when 0.400 moles of C6H12O6 are used?
The fermentation of glucose C6H12O6 produces ethanol, C2H5OH and CO2.The equation for the reaction is as follows: C6H12O6 --> 2C2H5OH + 2CO2.

How many grams of C2H5OH are produced when 0.400 moles of C6H12O6 are used?
1
Expert's answer
2017-10-03T14:55:07-0400
1) 2H2 + O2 = 2H2O
m(H2O)=n*Mr=(1/2)*2)*18=18 g.

2) 3H2+N2=2NH3
n(H2)=3/3=1 mol
n(N2=3/1= 3 mol
x=3-1= 2 mol unused

3) C6H12O6 --> 2C2H5OH + 2CO2
n(CO2)=(0,400/1)*2=0,800 mol
m(CO2)=n*Mr=0,800*48=35,2 g.
n(C2H5OH)=(0,400/1)*2=0,800 mol
m(C2H5OH)=0,800*46=36,8 g.